# DeFine: Enhancing LLM Decision-Making with Factor Profiles and Analogical Reasoning

### v1

@@ 9 Limitations
The effectiveness of the DeFine framework, as presented in this paper, is predominantly based on controlled experimental conditions. While the framework has been designed to enhance decision-making capabilities through the use of probabilistic factor profiles and analogical reasoning, actual outcomes may vary when applied in real-world scenarios. Users should be aware that the framework’s performance can be influenced by various external factors including data quality, context-specific nuances, and the dynamic nature of real-world environments. We encourage users to consider these variables when implementing and adapting the DeFine approach to ensure its optimal application and to mitigate potential discrepancies between expected and actual results.
Appendix A Appendix
[…]
PromptBox s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark, colback=light!30!white, fonttitle=, fontupper=, title=Constructing a Factor Profile from an Earnings Call Transcript,
{PromptBox}
System Message
@@ Appendix A Appendix
You are a financial analyst specializing in earnings call transcripts. You will receive the complete transcript of an earnings call, which includes both the prepared remarks and the Q&A session. Your job is to identify the key factors from the transcript and assign probabilities to the potential outcomes of these factors.
User Message
@@
Your task is to conduct a comprehensive analysis of the earnings call transcript below. Be sure to accurately capture the important factors and estimate the likelihood of each factor resulting in specific outcomes.
@@
Figure 5: Constructing a Factor Profile from an Earnings Call Transcript
[…]
histPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt to Analyze Trends Based on Historical Financial Metrics,
[…]
ARPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=Using Analogical Reasoning to Make Investment Decisions,
{ARPrompt}
System Message
You’re a financial analyst who specializes in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message
Here are several example company profiles. Each profile highlights key factors from an earnings call transcript and probabilities for potential outcomes based on those factors. Each profile represents a specific company and is based on its historical earnings call data. Your job is to pick the most analogous example and use its strategy to solve the initial problem.
@@
Figure 6: Using Analogical Reasoning to Make Investment Decisions
[…]
CoTPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt that Uses Chain-of-Thought to Make Investment Decisions,
{CoTPrompt}
System Message
You’re a financial analyst spcializing in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message
 - Action 1: strong buy: The stock price will increase by more than 5%
@@
Figure 7: A Prompt that Uses Chain-of-Thought to Make Investment Decisions
{histPrompt}
System Message
You are a financial analyst specializing in historical data analysis, including stock prices, earnings per share (EPS), and revenue. Your goal is to assess the likelihood of different market trends based on past data.
User Message
The potential outcomes to consider are: {bullish, stable, and bearish}. For each outcome, please assign a likelihood level from the following options: {very likely, likely, somewhat likely, somewhat unlikely, unlikely, very unlikely}.
@@
{… until the date of the earnings announcement.}
@@
Figure 8: A Prompt to Analyze Trends Based on Historical Financial Metrics

### v2

@@ 8 Conclusion
Acknowledgements
We are grateful to the reviewers for their insightful feedback, which has helped improve our paper. This research has been partially supported by the NSF CAREER award, #2303655.
@@ 9 Limitations
The DeFine framework, as detailed in our paper, offers a promising approach to enhancing decision-making through its innovative use of probabilistic factor profiles and analogical reasoning. Developed under carefully controlled experimental conditions, its potential is noteworthy. However, it’s important to acknowledge that its performance might vary in the complexity of real-world settings due to factors such as data quality and specific contextual nuances. We encourage users to consider these variables when adapting the DeFine approach to ensure its optimal application and to mitigate potential discrepancies between expected and actual results.
Appendix A Influential Factors
Transcript Excerpt from Delta Air Lines (DAL) Earnings Call [Prepared Remarks:] >> Operator Good morning, everyone, and welcome to the Delta Air Lines September-quarter 2021 financial results conference call. My name is Jen, and I will be your coordinator. [Operator instructions] As a reminder, todays call is being recorded. I would now like to turn the conference over to Ms. Julie Stewart, vice president of investor relations... >> Julie Stewart -- Vice President of Investor Relations Thank you, Jen. Good morning, everyone, and thanks for joining us for our September-quarter 2021 earnings call. Joining us from Atlanta today are CEO, Ed Bastian; our president, Glen Hauenstein; our CFO, Dan Janki. And Ed will open the call with an overview of Deltas performance and strategy. Glen will provide an update on the revenue environment and our brand momentum, and Dan will discuss cost fleet and our balance sheet. Similar to last quarters call, weve scheduled todays call for 90 minutes to make sure we have plenty of time for questions. [Operator instructions] After the analyst Q\&A, we will move to our media questions, after which, Ed will provide a brief closing statement. Todays discussion contains forward-looking statements that represent our beliefs or expectations about future events. All forward-looking statements involve risks and uncertainties that could cause the actual results to differ materially from the forward-looking statements. Some of the factors that may cause such differences are described in Deltas SEC filings. We also discuss non-GAAP financial measures, and all results exclude special items unless otherwise noted. You can find a reconciliation of our non-GAAP measures on the Investor Relations page at ir.delta.com. And with that, Ill turn the call over to Ed. >> Ed Bastian -- Chief Executive Officer Well, thank you, Julie, and good morning, everyone. Appreciate you joining us this morning. The September quarter marked another important milestone in our recovery. We achieved our first quarterly profit since the start of the pandemic with a pre-tax result of \$216 million and a pre-tax margin of nearly 3% despite still missing one-third of our revenue base compared to the same period in 2019... [omitted.] [Questions & Answers:] >> Operator Thank you. And well go first to Jamie Baker with J.P. Morgan. >> Jamie Baker -- J.P. Morgan -- Analyst Hey. Good morning, everybody. First question goes potentially to Glen and Dan. So pre-COVID, I had asked Paul about the amount of time that it would typically take Delta to recalibrate the higher fuel prices. Im not staring at the transcript, but his estimate at the time was four to six months, which was an improvement from historic levels. So my question, I guess, for Glen is whether the booking curve is steep enough right now that you might actually be able to recapture the top line more quickly than that. And similarly, for Dan, whether theres anything we should be taking on the cost or operations side that could accelerate the process. Im basically just trying to understand whether four to six months is still the right estimate for us to be using. >> Glen Hauenstein -- President Well, I would just comment, I think were a bit in uncharted territory here as the recovery continues. And while I think it might be difficult in the very short run, despite the fact that the booking curve has moved in a bit, that I would estimate that, that four to six months is about right because we believe that demand and capacity will fall back into a very good equilibrium by next spring which would put you inside that window...
Figure 5: Transcript Excerpt from Delta Air Lines (DAL) Earnings Call. Adapted from Lu et al. (2025).
@@ Appendix A Influential Factors
Building a Factor Profile from the Earnings Call Transcript System Message You are a financial analyst specializing in earnings call transcripts. You will receive the complete transcript of an earnings call, which includes both the prepared remarks and the Q&A session. Your job is to identify the key factors from the transcript and assign probabilities to the potential outcomes of these factors.
@@
User Message Your task is to conduct a comprehensive analysis of the earnings call transcript below. Be sure to accurately capture the important factors and estimate the likelihood of each factor resulting in specific outcomes.
@@
Figure 6: Building a Factor Profile from the Earnings Call Transcript
Applying Analogical Reasoning to Investment Decisions System Message You’re a financial analyst who specializes in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message Here are several example company profiles. Each profile highlights key factors from an earnings call transcript and probabilities for potential outcomes based on those factors. Each profile represents a specific company and is based on its historical earnings call data. Your job is to pick the most analogous example and use its strategy to solve the initial problem.
@@
Figure 7: Applying Analogical Reasoning to Investment Decisions.
Chain-of-Thought Prompt for Investment Decisions System Message You’re a financial analyst spcializing in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message Based on your analysis of the earnings call for {Company Name} held on {Announcement Date}, decide on the most likely analyst recommendation for the next 30 days from these options:
- Action 1: strong buy: The stock price will increase by more than 5%
@@
Figure 8: Chain-of-Thought Prompt for Soliciting Investment Decisions.
Prompt for Analyzing Trends in Historical Financial Data System Message You are a financial analyst specializing in historical data analysis, including stock prices, earnings per share (EPS), and revenue. Your goal is to assess the likelihood of different market trends based on past data.
User Message The potential outcomes to consider are: {bullish, stable, and bearish}. For each outcome, please assign a likelihood level from the following options: {very likely, likely, somewhat likely, somewhat unlikely, unlikely, very unlikely}.
@@
{...until the date of the earnings announcement.}
@@
Figure 9: Prompt for Analyzing Trends in Historical Financial Data.